OpenStudy (anonymous):

Having a brainfreeze... (x-2/3)^2 =14 Can someone help me with this? I would be so grateful :)

OpenStudy (bibby):

the inverse of raising a quantity to a power is taking the same root of it. for example: \((\sqrt{x})^2=x\\(\sqrt[3]{87})^3=87\)

OpenStudy (anonymous):

so it would just be x-2/3=14? would I take the square root of 14, or just solve it?

OpenStudy (bibby):

take the square root of both sides to get rid of the ^2, yeah

OpenStudy (anonymous):

Oh! ok, I remember now... Thank you so much!

OpenStudy (bibby):

feel free to post your answer here if you want it checked anytime~

OpenStudy (anonymous):

So I got x=sqroot14+ 2/3 is that right? Or is this one of those problems where I need to put a +_ in front?

OpenStudy (bibby):

\((\dfrac{x-2}{3})^2 =14\\\sqrt{(\dfrac{x-2}{3})^\cancel 2} =\pm\sqrt{14}\\\dfrac{x-2}{3}=\pm\sqrt{14}\\x-2= \pm 3\sqrt{14}\\x= 2+\pm 3\sqrt{14}\implies x=2+3\sqrt{14}\\x=2-3\sqrt{14}\)

OpenStudy (anonymous):

Okay awesome. I think I got it. Thanks for explaining!

OpenStudy (bibby):

anytime btw you can also do it by squaring out the terms individually. recall that a/b^2=a^2/b^2

OpenStudy (bibby):

and with that you can get http://www.wolframalpha.com/input/?i=%28x%5E2-4x%2B4%29%3D14*9 (x^2-4x+4)/9=14 (x^2-4x+4)=126 then solve that

OpenStudy (anonymous):

Okay sweet! This has really helped a ton!